how many atoms are in 5.081 moles of Zn? In 1 lb. box of salt (NaCl), 1 lb=454 gram, how many atoms of Cl are in the box?
1
Expert's answer
2017-04-11T01:47:08-0400
1 mole of substance contains 6.022 · 1023 parts (Avogadro’s number).
5.081 moles of Zn contain atoms: 5.081 · 6.022 · 1023 = 30.598 · 1023

M (NaCl) = 58.442 g/mol
n (NaCl) = m/M = 454/58.442 = 7.768 mol
N = 7.768 · 6.022 · 1023 = 46.781 · 1023